OpenStudy (melissa_something):

Why is the other vertical asymptote 0?

OpenStudy (melissa_something):

Idk if this matters but the question is x+3 / x(x-4)

OpenStudy (phi):

***Idk if this matters but the question is x+3 / x(x-4)*** yes , it matters; otherwise no one could answer the question if you have a fraction like for example \[ \frac{3}{0.001}\] that simplifies to 3000 (if you multiply top and bottom by 1000, you will get that answer, or if you use a calculator you will get 3000) if we had an even smaller number, like \[ \frac{3}{0.000001}\] that simplifies to 3000000, even bigger. The point is: dividing by a *tiny* number gives a large number for the answer now look at \[ \frac{x+3}{x(x-4)} \] as x "gets near" zero, as for example: 0.0001 or 0.0000001 or 0.000000001 the answer will get huge. It "asymptotes" btw, the same thing happens when x gets near 4 because x-4 gets close to zero, and the same thing happens